SpaceRef

SpaceRef


NOAA GOES SATOPS Morning Report 21 May 2001

Status Report From: China National Space Administration
Posted: Monday, May 21, 2001

SOCC Relocation Update: The DMSP OPS are move will occur tomorrow, May 22, starting at 09:00am in order to have it done by the launch simulation being held the first week of June. The POES OPS area move will occur the week of May 28th.

POES

No change in the status of the POES spacecraft.

NOAA-14 rev 32895 / F at 1418z on May 18: GAC1B received noisy from spacecraft with 94.9% of the expected data recovered. Another recovery attempt was not possible due to spacecraft recorder scheduling limitations.

The NOAA-11 SBUV operating range will be modified today via flex memory load to measure wavelengths in the 306-340 nm range. Recently, due to a degrading grating drive mechanism, the NOAA-11 SBUV instrument was limited to only providing total ozone data. The wavelengths are currently confined to a range of 313 - 340nm, instead of the normal stepping profile over a much larger range. The requested 306 - 340nm range will slightly extend the operating range of the SBUV grating drive, which will provide the ozone users additional valuable ozone information and facilitate instrument-to-instrument comparisons studies with NASA's EP TOMS instrument. The extended range will operate the instrument in a known problem-free area.

POES operations were nominal over the past 72 hours.

DMSP

No change in the status of the DMSP spacecraft.

The multiple data losses continued throughout the weekend as a result of the on-going problems at the POGO site. The DMSP operations crews minimized data losses via extensive site call-ups and data chasing. As of 0444z on May 21, a ground system cable replacement at POGO seems to have recovered site capability. The site is not yet declared Green, but SOCC will use site as scheduled. Since 0444z, performance appears nominal.

DMSP operations were nominal over the past 72 hours.

GOES

No change in the status of the GOES spacecraft.

At 0024z on May 19, the GOES-10 schedule suspended due to encryptor 'B' failure; lost 0024Z sounding as a result.

On May 19 at 0745z, 0754z, 0805z, 0817, 0821 and 0847z, GOES-10 frame breaks were observed and attributed to high BER due to local thunderstorms at CDA.

On May 19 at 0821z, 0832z, 0840z and 0848z, GOES-8 frame breaks were observed and attributed to high BER due to local thunderstorms at CDA.

At 1427z on May 19, a GOES-10 frame break was observed and attributed to loss of signal synchronization.

GOES-11 North/South Maneuver and Normal On-Orbit Mode Testing Activities: The return of GOES-11 to normal on-orbit mode will run from May 29 to June 11. During this time a North/South maneuver and an East/West maneuver will be done, the GOES-12 GVAR software will be tested by tape transmissions, the GOES-12 Solar X-Ray Imager (SXI) MDL data reception will be tested by tape transmissions, the SAR and DCPR transponders will be recalibrated, and the GOES-11 imager and sounder will be operated for one week. GOES-11 will then do a re-spin to return to storage mode.

GOES operations were nominal over the past 72 hours.

Questions or comments to Tim.Walsh@noaa.gov

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.